A nurse wants to prepare a 1.5 % (m/v) silver nitrite solution from 24 mL of 8 % (m/v) stock solution of sliver nitrate. How much water should be added to the 24 ml of the stock solution? Please show your work.
Answer #1

Answer:

Given data

A nurse want to prepare a 1.5% (m/v) silver nitrate solution from 24 ml of 8% (m/v) stock solution of silver nitrate

CV = CV
(1.5)(V) =(8)(24)
V = 128 mL
Volume to add is 128 - 24 = 104 mL....

Obviously this. all assumes volumes are added substance and, strickly, this is not genuine.

The main thing to perceive however is to guarantee the last volume is 128....but adjusting to the significant of your concern,

130 would be fine.